OpenStudy (ggray):

Please help, will medal Which is equal to 42 • 48? A: 416 B: 410 C:406 D:4-6

OpenStudy (anonymous):

ummmm?

Nnesha (nnesha):

please use `^` for exponents \[\huge\rm 4^2 \times 4^8\] this is ur question same with options

OpenStudy (anonymous):

What do you think it is? :)

Nnesha (nnesha):

when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\] apply the exponent rule

OpenStudy (ggray):

i think it is d

Nnesha (nnesha):

why do you think its D ?

OpenStudy (ggray):

well if you subtract the exponents you get 2-8=-6

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\] apply the exponent rule \(\color{blue}{\text{End of Quote}}\)

OpenStudy (ggray):

oh okay thank you!

Nnesha (nnesha):

alright what did yo uget ?? "

OpenStudy (ggray):

b?

Nnesha (nnesha):

yep 2+8 =10

OpenStudy (ggray):

thanks

OpenStudy (ggray):

Can you help me on another one real fast?